a generator at a utility company produced 100 A of current at 4000 V. the voltage is stepped up to 24oooo V by a transformer before it is sent on a high transmission voltage line. what is the current in the transmission line

Grade:12th pass

2 Answers

Piyush Maheshwari
6 years ago
In stepping up, Power remains same.
VI= const.
100 * 4000 = i * 240000
Hence current of stepped up supply i=10/6 = 1.66 A
SR Roy
6 years ago
 
 
 
In stepping up, Power remains same.
VI= const.
100 * 4000 = i * 240000
Hence current of stepped up supply i=10/6 = 1.66 A